Create equations in two or more variables to represent relationships between quantities.
10. Isaac is getting his own cell phone plan and wants to make the best decision. He decides to compare cell phone plans. MobileONE has an unlimited talk plan for a flat fee of $80/month. Talk-A-Lot charges $0.15 per minute with a monthly fee of $25. By-The-Minute has no monthly fee, but is $0.50 per minute. Graph these cell phone plans. Which plan is cheapest if Isaac is using approximately 50 minutes each month? Which plan is cheapest if Isaac is using approximately 100 minutes each month? Which plan is cheapest if Isaac is using approximately 400 minutes each month?

Answer by addingup(3677)

MobileONE: 80+0x = 80/month
Talk-A-Lot: 25+0.15x =
By-The-Minute: 0+0.50x =
Substitute the x with number of minutes and use your calculator to get your answers.